OpenStudy (anonymous):

Can someone please help me with part 2 and 3? Part 1: Explain, in complete sentences, whether or not the polynomial x2 – 16x + 64 fits one of the special patterns. If so, which one? (2 points) binomial square Part 2: Explain how the polynomial is factored. (2 points) Part 3: Provide the factors. (2 points)

OpenStudy (anonymous):

@Ashleyisakitty @april115 @AnElephant @Algebra2013 @ashley_f97 @alivia97 @breton @Bea_26 @baby456 @blah124 @banana9715 @bashirk @christianfilms11 @churrity @chaylacryder @caitlynmcdaniel @coolgirl3517 @captainmico88 @dan815 @DebbieG @dape @dirtydan667 @dlapointe25 @danielbarrios @ExplainItLikeImFive @Falco276 @Firejay5 @ganeshie8 @geoffb @greenfrostgiant @hereweg0 @heyitzkellymarie1484 @heberv @iceicebaby @Its_Sonic @izhangy @its_martahuo @jim_thompson5910 @Jamierox4ev3r @jagatuba @JackJones @Joshyface @JetWave600 @kelliegirl33 @KingGeorge @kym02 @kchill @Luigi0210 @Loser66 @Loujoelou @love_jessika15 @litchlani @lavalava @mathstudent55 @mathslover @myko @miadeex @marylou004 @music101 @nayeaddo92 @oldrin.bataku @orple8 @ovanbc @phi @Preetha @pgpilot326 @paiger16 @robtobey @Ryaan @RoseDryer @rasperrystumper @Requiem @raaju @satellite73 @selrachcw95 @sassoue @StevenGlendasBoy @sierrrrasellsssells @Samanthaa_Ann13 @thomaster @timo86m @ttop0816 @Tayloresc @tshark14 @iceicebaby @Its_Sonic @izhangy @its_martahuo @UsaChan<3 @valiyuh @Xtasy0xy @xxImHisGirlxx @yummydum @ybarrap @yeshlii @YolandaL @zzr0ck3r

OpenStudy (anonymous):

every one on openstudy ever ^

OpenStudy (april115):

Damn

OpenStudy (anonymous):

U_U

OpenStudy (anonymous):

what a crowd..lol

OpenStudy (luigi0210):

Hey Everybody, Welcome to Openstudy

OpenStudy (anonymous):

I hate math

OpenStudy (anonymous):

Thanks @Luigi0210 you're the first one who welcomed me!

OpenStudy (anonymous):

your polynomial is (x - 8)^2...what pattern is that ?

OpenStudy (anonymous):

@christianfilms11 @churrity @chaylacryder @caitlynmcdaniel @coolgirl3517 @captainmico88 @dan815 @DebbieG @dape @dirtydan667 @dlapointe25 @danielbarrios @ExplainItLikeImFive @Falco276 @Firejay5 @ganeshie8 @geoffb @greenfrostgiant @hereweg0 @heyitzkellymarie1484 @heberv @iceicebaby @Its_Sonic @izhangy @its_martahuo @jim_thompson5910 @Jamierox4ev3r @jagatuba @JackJones @Joshyface @JetWave600 @kelliegirl33 @KingGeorge @kym02 @kchill @Luigi0210 @Loser66 @Loujoelou @love_jessika15 @litchlani @lavalava @mathstudent55 @mathslover @myko @miadeex @marylou004 @music101 @nayeaddo92 @oldrin.bataku @orple8 @ovanbc @phi @Preetha @pgpilot326 @paiger16 @robtobey @Ryaan @RoseDryer @rasperrystumper @Requiem @raaju @satellite73 @selrachcw95 @sassoue @StevenGlendasBoy @sierrrrasellsssells @Samanthaa_Ann13 @thomaster @timo86m @ttop0816 @Tayloresc @tshark14 @iceicebaby @Its_Sonic @izhangy @its_martahuo @UsaChan<3 @valiyuh @Xtasy0xy @xxImHisGirlxx @yummydum @ybarrap @yeshlii @YolandaL @zzr0ck3r

OpenStudy (anonymous):

haha

OpenStudy (anonymous):

(x-8)(x-8)=0 is the factoring of the problem and the factor is 8

OpenStudy (anonymous):

:p

OpenStudy (luigi0210):

Dan, I'm right here

OpenStudy (anonymous):

@paiger16 no its not

OpenStudy (april115):

Luigi do you see these yellows and crap colors and pale oranges i think were already welcomed! A LONG TIME AGO >.>

OpenStudy (anonymous):

(x - 8)^2

OpenStudy (anonymous):

I need help with Part 2 and Part 2. Part 1 is squared binomial.

OpenStudy (anonymous):

nope. that would end up being x^2-64

OpenStudy (anonymous):

hold on 4 a sec

OpenStudy (anonymous):

i dont know what do you think

OpenStudy (anonymous):

just kidding nvm

OpenStudy (anonymous):

No, Seriously! no one welcomed me when I came here for the first time! @april115

OpenStudy (anonymous):

part 2: two numbers v and w multiply to give 64 and sum to give -16 part 3: You have x^2+bx+c where (1) b=v+w (2) c=v*w Two equations with two unknowns that you can easily solve.

OpenStudy (anonymous):

i put all those names in a word document to have them for later ....lol

OpenStudy (anonymous):

wow.. a lot of people here..

OpenStudy (april115):

LMAO my bad

OpenStudy (dan815):